# Evaluating AI systems under uncertain ground truth: a case study in dermatology

### v1

@@ Abstract
For safety, AI systems in health undergo thorough evaluations before deployment, validating their predictions against a so-called ground truth. Importantly, this ground truth is assumed known and fixed, i.e., certain. However, especially in health settings, this is actually not the case and the ground truth may be uncertain. Unfortunately, this is largely ignored in standard evaluation of AI models but can have severe consequences such as overestimating the future performance (Gordon et al., 2021). To avoid this, we measure the effects of ground truth uncertainty, which we assume decomposes into two main components: annotation uncertainty which stems from the lack of reliable annotations, and inherent uncertainty due to limited observational information. This ground truth uncertainty is ignored when estimating the ground truth by deterministically aggregating annotations, e.g., by majority voting or averaging. In contrast, we propose a framework where aggregation is done using a statistical model. Specifically, we frame aggregation of annotations as posterior inference of so-called plausibilities, representing distributions over classes in a classification setting, subject to a hyper-parameter encoding annotator reliability. Based on this model, we propose a metric for measuring annotation uncertainty and provide uncertainty-adjusted metrics for performance evaluation. We present a case study applying our framework to skin condition classification from images (Liu et al., 2020) where annotations are provided in the form of differential diagnoses, modeled as partial ranking of conditions. The deterministic adjudication process called inverse rank normalization (IRN) from previous work ignores ground truth uncertainty in evaluation. Instead, we present two alternative statistical models: a probabilistic version of IRN and a Plackett–Luce-based model (Plackett, 1975; Luce, 2012). We find that a large portion of the dataset exhibits significant ground truth uncertainty and standard IRN-based evaluation severely over-estimates performance without providing uncertainty estimates. In contrast, our framework provides uncertainty estimates on common metrics of interest such as top-$k$ accuracy and average overlap, showing that performance can change multiple percentage points depending on the annotator reliability which directly impacts model selection.
@@ 1 Introduction
Prior to usage, predictive AI models in health are usually evaluated by comparing model predictions on a held-out test set annotated with a corresponding known ground truth. In a supervised classification context, this typically assumes the availability of a single, unique and certain ground truth per example. In almost all benchmarks for supervised learning, this ground truth is derived by deterministic aggregation of multiple human annotations, e.g., using simple majority voting or averaging, often ignoring any uncertainty or disagreement. For example, in medical diagnosis, doctors hired for labeling examples have to make complex decisions with very limited information. Where doctors would usually ask patients questions or perform additional tests, they instead have to come up with a working hypothesis of possible conditions, a so-called differential diagnosis. On ambiguous cases, doctors with different levels of experience in medicine, different expertise and biases, constrained by an imperfect labeling tool, will come to different conclusions (Freeman et al., 2021).
Instead of ignoring this disagreement and pretending the ground truth to be fixed and certain, we argue that it is important to acknowledge and address this disagreement. This is to ensure the best possible outcome for the patient because disagreement often reflects individual challenges that when ignored pose significant risks. Specifically, we state that ground truth is uncertain. This uncertainty can be decomposed into annotation uncertainty and inherent uncertainty. The former stems from an imperfect labeling process: even expert annotators can make mistakes; tasks might be subjective, annotators might be biased, inexperienced in using the labeling tool or simply lack experience in the labeling task. The inherent uncertainty, on the other hand, stems from limited observational information. In the example above, there might be ambiguous cases where inferring a clear condition for a patient solely on a single image or a short report, can be extremely difficult. This also includes what related work calls task ambiguity (Uma et al., 2022) (e.g., if classes are not agreed upon (Phene et al., 2019; Medeiros et al., 2023)). In practice, ground truth uncertainty is manifested through disagreement among annotators (e.g., measured by inter-annotator disagreement) and it is challenging to attribute disagreement to annotation or inherent uncertainty (Abercrombie et al., 2023; Röttger et al., 2022). However, while recruiting a larger group of annotators, training them better or improving the labeling tool can reduce annotation uncertainty (usually at a very high cost), inherent uncertainty is generally irresolvable (Schaekermann et al., 2016). This also holds for adjudication (Schaekermann et al., 2019b; Duggan et al., 2021). Moreover, deterministically aggregating more annotators may eliminate minority views (Field et al., 2021).
@@
Figure 1: Ground truth uncertainty in supervised AI systems such as classification: The unobserved ground truth $y^{*}$ is assumed to yield an observation $x$. This observation can contain limited information giving rise to inherent uncertainty where the true label might not be clearly identifiable purely from $x$. Nevertheless, experts are asked to provide annotations $b^{1},\dots,b^{R}$ based on $x$. These are subject to annotation uncertainty caused by the lack of reliable annotations. See Figure 5 for a concrete example in skin condition classification. These annotations are then typically (deterministically) aggregated and a single estimate $y$ of the true label $y^{*}$ is extracted. Evaluation of AI models uses the estimate $y$ as ground truth, ignoring any ground truth uncertainty. In contrast, we use statistical aggregation to explicitly model distributions over ground truth labels, so-called plausibilities $\lambda$, representing categorical distributions over classes. The distribution over $\lambda$ captures annotation uncertainty; the plausibilities themselves may capture inherent uncertainty. We discuss how to measure annotation uncertainty and evaluate AI systems on top of these plausibilities.
@@
We can observe annotator disagreement in many standard datasets (see Figure 3 for an example on CIFAR10 (Krizhevsky, 2009; Peterson et al., 2019)). For many of these benchmarks, ground truth uncertainty is limited to a small fraction of examples while the ground truth of the majority of examples can be trusted. In medicine, in contrast, it is common that a significant portion of examples is subject to ground truth uncertainty, observed by high annotator disagreement (Schaekermann et al., 2016). For example, this has recently been shown in skin condition classification (Jain et al., 2021; Eng et al., 2019), but also holds beyond health, e.g., in toxicity classification (see Figure 9). As a result, the problem of ground truth uncertainty has been recognized in several previous works; see e.g. (Sculley, 2007; Cabitza et al., 2020; Northcutt et al., 2021; Uma et al., 2021; Gordon et al., 2021; Davani et al., 2022; Plank, 2022; Leonardelli et al., 2023). Often, however, the focus is on mitigating the symptoms of ground truth uncertainty rather than tackling it directly. For example, there is a large body of work on dealing with label noise (Northcutt et al., 2021). In contrast, we believe that many instances of label errors stem from ignoring the underlying disagreement. In cases where ground truth uncertainty is modeled explicitly, related work focuses on training (Welinder et al., 2010; Rodrigues and Pereira, 2018; Guan et al., 2018) but still assumes certain ground truth for evaluation. This is despite Maier-Hein et al. (2018) explicitly highlighting the impact of ignoring annotator disagreement in evaluation on model selection and ranking. As illustrated in (Gordon et al., 2021), ignoring disagreement by deterministically aggregating annotations may lead to misleading and fragile results when assessing the future performance of AI systems.
We propose here a general framework for measuring and evaluating with ground truth uncertainty based on statistical aggregation of annotations. Essentially, we pose aggregation as a posterior inference task following Figure 1: given multiple annotations and a prior knowledge of annotator reliability, we infer so-called plausibilities. In a classification setting, these plausibilities represent categorical distributions over classes. The variation in plausibilities captures annotation uncertainty. The plausibilities themselves capture inherent uncertainty depending on the entropy of the corresponding categorical distributions. Reliability can be thought of as a model parameter that provides a prior over the expected variation in plausibilities (cf. Figure 2). It could be informed by domain experts or tuned on data (similar to work in crowd sourcing (Yan et al., 2014; Zheng et al., 2017)). However, as quantifying annotator reliability is difficult, we assume it to be a free parameter during evaluation. Then, we propose a measure of annotation certainty to quantify the uncertainty of any label being the true ground truth. This allows us to quantify annotation uncertainty on individual examples as well as on whole datasets. In addition, we present uncertainty-adjusted variants of common classification metrics such as top-k accuracy or average overlap. Altogether, we provide a comprehensive strategy to evaluate AI systems for health under ground truth uncertainty.
In this paper, we apply our framework to skin condition classification from images in dermatology following the setting of (Liu et al., 2020). Here, annotations are expressed as differential diagnoses, i.e., partial rankings. Because classifying skin conditions purely from images is an incredibly difficult task, there is significant disagreement among annotations for a large portion of the dataset. Following (Liu et al., 2020), we consider inverse rank normalization (IRN) as baseline deterministic aggregation method. IRN can be thought of as providing a plausibility point estimate and prior work typically used the top-1 label as ground truth for evaluation – ignoring both annotation and inherent uncertainty. Instead, we propose two alternative statistical models: a probabilistic interpretation of IRN and a Plackett–Luce-based model (Plackett, 1975; Luce, 2012) specifically adapted to partial rankings. Our experiments highlight the high annotation uncertainty in this setting. Moreover, they show that previous IRN-based evaluation significantly over-estimates classifier performance and disregards large variations in performance due to ground truth uncertainty, hindering model selection. We discussed these results with dermatologists and highlight medical implications such as the inability to clearly categorize cases by risk.
@@
2 Framework for evaluation under uncertain ground truth
@@
This section introduces our framework for evaluation of AI systems with ground truth uncertainty. While this paper focuses on a case study in dermatology, we present our framework in general terms and discuss technical details of applying it to the differential diagnosis annotations from (Liu et al., 2020) in Appendix B. We start by introducing notation and give an intuition of our approach on a toy example as well as a standard image recognition task, namely CIFAR10 (Krizhevsky, 2009). We then formalize this intuition by presenting the statistical model we use for modeling how annotator opinions are aggregated. Based on this model, we present measures for annotation uncertainty as well as uncertainty adjusted performance metrics for evaluating AI models.
@@ 2.1 Notation and introductory examples
Figure 2: Illustration of annotation certainty on a toy dataset. Left: Synthetic dataset showing one-dimensional observations $x$ and corresponding true distribution $p(y|x)$ over labels $y$ for three imaginary labels Benign1, Benign2 and Cancer. Example (ii) is particularly ambiguous in the sense that $p(y|x)$ is not crisp. Middle left: For illustration, we plot the distribution $p(y|x)$ on the 3-simplex. Middle and middle right: Modeling aggregation of annotator opinions $b$ statistically, see text, allows us to sample plausibilities $\lambda^{(m)}\sim p(\lambda|b,x)$. The spread of these plausibilities around the actual true distributions $p(y|x)$ captures the annotation uncertainty. This is influenced by the annotations as well as a prior reliability parameter $\gamma$. This parameter reflects our a prior trust in the annotators. Right: Measuring how often the top-1 label changes given plausibility samples $\lambda^{(m)}$, we can compute annotation certainty for all examples. We plot annotation certainty across different reliabilities $\gamma$ (black to gray) indicating the effect of different prior trust levels.
@@
For illustration and introducing notation, we consider the synthetic toy dataset of Figure 2 (left). Here, observations $x$ are illustrated as one-dimensional on the x-axis and we plot the true distribution $p(y|x)$ for three different classes (blue, orange and violet) on the y-axis. We highlight three examples, (i), (ii) and (iii), where the latter two are inherently ambiguous, i.e., the corresponding distribution $p(y|x)$ is not crisp. This is intuitively visualized when plotting the distributions $p(y|x)$ on a 3-simplex where the corners would correspond to crisp distributions (middle left). Of course, we never observe the true $p(y|x)$ for these points. Instead, we assume access to a finite set of expert annotations $b$. For simplicity, in this example we assume these annotations to be single labels (sampled from the true $p(y|x)$). Then, we aggregate these opinions to obtain an approximation $\hat{\lambda}$ of $p(y|x)$ and select a label to use as ground truth based on, e.g., majority voting. This represents how labels for many common benchmarks such as CIFAR, ImageNet (Russakovsky et al., 2015) or Hateful Memes (Kiela et al., 2020) have been obtained. We refer to $\hat{\lambda}$ as plausibilities as they construct a distribution $p(y|\hat{\lambda})$ over labels from which the $\arg\max$ corresponds to the majority voted label. However, this approach ignores any uncertainty present in the annotations $b$.
@@
Instead, we assume a distribution $p(\lambda|b,x)$ over plausibilities. Here, we use a Dirichlet distribution with concentration parameters reflecting the annotator opinions as well as a prior reliability parameter ($\gamma$ in Figure 2; also see Appendix A for details). This reliability parameter will quantify our a priori trust in the annotators. For example, we expect reliability to increase with the number of annotators or their expertise and training. However, choosing or estimating reliability can be challenging or require additional information. Instead, we take the reliability to be a free parameter, allowing us different views on the data [^1] We use a single global reliability parameter across all annotators for simplicity and evaluation in Section 3.. This is illustrated in Figure 2 (middle and middle right), showing plausibilities $\lambda^{(m)}\sim p(\lambda|b,x)$, $m\in[M]$, sampled from the statistical aggregation model on the 3-simplex. The spread in these plausibilities represents annotation uncertainty and can be reduced using a higher reliability. The position on the simplex represents inherent uncertainty which is lower for plausibilities close to the corners such as (i) and higher for plausibilities close to the center such as (ii). Then, we compute the top-1 label for every sampled $\lambda^{(m)}$, i.e., the label with the largest plausibility $\arg\max_{k\in[K]}\lambda^{(m)}_{k}$. This allows us to measure annotation certainty (defined formally in Section 2.3) – if the top-1 label is always the same label, there is no annotation uncertainty. Specifically we can measure the fraction of plausibilities $\lambda^{(m)}$ among all $M$ samples where the top-1 label is $k$. Taking the maximum over all labels $k$ defines annotation certainty. In Figure 2 (middle) we plot this annotation certainty for different reliabilities (black and gray) and we clearly see that annotation certainty is consistently low for example (ii), while for (iii) annotation certainty increases with higher reliability. This also allows us to easily summarize annotation uncertainty across the whole dataset.
Figure 3: We plot annotation certainty and uncertainty-adjusted (UA) accuracy for CIFAR10-H (Krizhevsky, 2009; Peterson et al., 2019), highlighting the 1000 examples with lowest annotation certainty. There is a small set of $178$ images with certainty below $99\%$ for which annotators tend to disagree. Given the large number of annotators ($50$), this likely indicates examples with inherent uncertainty, as shown for two examples on the right (images and histogram of annotations).
To move from this toy example to a real benchmark, we consider CIFAR10 (Krizhevsky, 2009) with the human annotations from (Peterson et al., 2019). As above, each annotator provides a single label such that we can apply the same methodology. Figure 3 (left) reports the corresponding annotation certainty alongside an uncertainty-adjusted version of accuracy. As can be seen, annotators agree for the majority of examples, resulting in certainties close to $1$. Nevertheless, on roughly $178$ test examples, certainty is below $0.99$. As there are 50 annotators per example with only 10 classes, it is likely that these are actually difficult cases, i.e. cases with inherent uncertainty, as confirmed in Figure 3. This corresponds to roughly $0.2\%$ of the examples. Interestingly, recent improvements in accuracy on CIFAR10 are often smaller than $0.2\%$ [^2] According to Papers with Code, https://paperswithcode.com/sota/image-classification-on-cifar-10.. To measure accuracy while taking annotation uncertainty into account, we measure how often the original CIFAR10 ground truth labels coincide with the top-1 labels obtained from sampled plausibilities. We call this uncertainty-adjusted accuracy (formally defined in Section 2.4) and, unsurprisingly, observe that the original labels of (Krizhevsky, 2009) (when interpreted as predictions) often perform poorly for examples with high annotation uncertainty (red in Figure 3). We expect the impact of ground truth uncertainty to be significantly more pronounced in settings with more disagreement, especially in our dermatology case study (Liu et al., 2020), but also in other domains. For example, Appendix A includes an example on Wikipedia Toxicity (Wulczyn et al., 2017) and we refer to (Leonardelli et al., 2021) for more examples in NLP.

### v2

# Evaluating medical AI systems in dermatology under uncertain ground truth

@@ Abstract
For safety, medical AI systems undergo thorough evaluations before deployment, validating their predictions against a ground truth which is assumed to be fixed and certain. However, in medical applications, this ground truth is often curated in the form of differential diagnoses provided by multiple experts. While a single differential diagnosis reflects the uncertainty in one expert assessment, multiple experts introduce another layer of uncertainty through potential disagreement. Both forms of uncertainty are ignored in standard evaluation which aggregates these differential diagnoses to a single label. In this paper, we show that ignoring uncertainty leads to overly optimistic estimates of model performance, therefore underestimating risk associated with particular diagnostic decisions. Moreover, point estimates largely ignore dramatic differences in uncertainty of individual cases. To this end, we propose a statistical aggregation approach, where we infer a distribution on probabilities of underlying medical condition candidates themselves, based on observed annotations. This formulation naturally accounts for the potential disagreements between different experts, as well as uncertainty stemming from individual differential diagnoses, capturing the entire ground truth uncertainty. Practically, our approach boils down to generating multiple samples of medical condition probabilities, then evaluating and averaging performance metrics based on these sampled probabilities, instead of relying on a single point estimate. This allows us to provide uncertainty-adjusted estimates of common metrics of interest such as top-$k$ accuracy and average overlap. In the skin condition classification problem of [58], our methodology reveals significant ground truth uncertainty for most data points and demonstrates that standard evaluation techniques can overestimate performance by several percentage points. We conclude that, while assuming a crisp ground truth may be acceptable for many AI applications, a more nuanced evaluation protocol acknowledging the inherent complexity and variability of differential diagnoses should be utilized in medical diagnosis.
keywords:
Keywords:
Dermatology
Evaluation
Uncertainty
Annotator disagreement
Label noise
journal: Medical Image Analysis
@@ 1 Introduction
Prior to deployment, predictive AI models are usually evaluated by comparing model predictions to a known ground truth on a held-out test set. In a supervised classification context, this typically assumes the availability of a unique and certain ground truth per example. Commonly used supervised training and evaluation procedures are usually built on this assumption. However, in certain domains, predominantly in health, such ground truth is often not available.
In medical diagnosis, uncertainty is the norm as doctors have to constantly make complex decisions with limited information. The common strategy to deal systematically with uncertainty is to come up with a differential diagnosis. A differential diagnosis is a working list of medical conditions ordered according to the perceived subjective plausibility of each condition. In a standard diagnosis process, a doctor would usually proceed to ask the patient further questions or perform additional tests to converge on a final diagnosis. However, in machine learning tasks annotated by such experts, their feedback remain in the form of differential diagnoses, reflecting the existing uncertainty in their assessment limited by the information provided. Moreover, multiple physicians with different backgrounds and medical experiences, constrained by an imperfect labeling tool, will typically provide different differentials, especially on ambiguous cases [38].
In medical AI applications, ground truth labels are usually derived by simplistic aggregations of multiple differential diagnoses to fit existing standard machine learning evaluation metrics, such as top-$k$ accuracy. The choice of this aggregation mechanism is often not made explicit in evaluations. More importantly, this largely ignores the uncertainty and wealth of information in the potentially disagreeing expert annotations. In this paper, we show that ignoring this uncertainty in expert annotations leads to overly confident and optimistic estimates of model performance, thereby underestimating risk associated with particular diagnostic decisions. Instead, we propose a principled evaluation procedure to obtain uncertainty-adjusted estimates of the real model performance.
1.1 Ground truth uncertainty and annotator disagreement
Throughout the paper, we consider two sources of ground truth uncertainty, annotation uncertainty and inherent uncertainty. Annotation uncertainty stems from labeling process complexities: even expert annotators can make mistakes, tasks might be subjective, annotators could be biased, or lack experience with the labeling tool or task. Inherent uncertainty arises from limited observational information, such as diagnosing a medical condition using a single medical image in the skin classification application explored here [58]. We ignore other sources of uncertainty like task ambiguity [105] that arises for example when individual classes are not clearly agreed upon [74, 64].
Ground truth uncertainty is observed most prominently through annotator disagreement. While inter-annotator disagreement can be measured through disagreement metrics, attributing it to annotation or inherent uncertainty remains challenging [1, 83]. Annotation uncertainty can potentially be reduced by expanding annotator groups, improving training, or refining labeling tools. In contrast, inherent uncertainty is generally irresolvable [89], even with costly adjudication protocols [91, 28]. Moreover, simple aggregations of opinions, such as majority voting, may eliminate minority views [35].
@@ 1.1 Ground truth uncertainty and annotator disagreement
Fig. 1: Ground truth uncertainty in supervised AI systems such as classification. The unobserved ground truth $y^{*}$ is assumed to yield an observation $x$. Here, $y^{*}$ could be the true underlying medical condition and the observation $x$ can be a medical image. Typically, $y^{*}$ is not uniquely identifiable from $x$ due to inherent uncertainty. To get an estimate of $y^{*}$, experts are asked to provide their annotations $b$ about a case. The experts can have different opinions how to interpret a case; therefore, to deal with annotation uncertainty instead of a single expert, multiple experts $R$ are consulted to give annotations $b^{1},\dots,b^{R}$ based on $x$. These annotations are then typically (deterministically) aggregated and a single estimate $y$ of the true label $y^{*}$ is extracted and evaluation proceeds, ignoring annotation or inherent uncertainty. See Figure 5 for a concrete example in skin condition classification. In contrast, we use statistical aggregation to explicitly model distributions over probabilities of ground truth labels, that we call plausibilities $\lambda$. The distribution over $\lambda$ captures annotation uncertainty; and a single plausibility vector $\lambda$ captures the inherent uncertainty of a case. We discuss how to evaluate AI systems using this modeling approach.
@@
While annotator disagreement can be observed in many standard datasets [52, 72], it is particularly common in medicine. For example, in skin condition classification [47, 30], a significant portion of examples is typically subject to ground truth uncertainty [89]. As a result, the problem of ground truth uncertainty has been recognized in several previous works; see e.g. [94, 14, 69, 104, 41, 23, 76, 56]. Often, however, the focus is on mitigating the symptoms of ground truth uncertainty rather than tackling it directly. For example, there is a large body of work on dealing with label noise [69]. In cases where ground truth uncertainty is modeled explicitly, related work focuses on training [113, 81, 43] but still assumes ground truth labels for evaluation. These shortcomings have recently been highlighted in [62, 41], but remain unaddressed.
1.2 Evaluation under ground truth uncertainty
To address evaluation under ground truth uncertainty, we propose a framework that considers a distribution over the ground truth rather than a single point estimate. This approach more faithfully captures the information communicated by annotators and is suitable for a broad spectrum of performance metrics beyond accuracy. Concretely, as illustrated in Figure 1, given multiple annotations $b^{1},\ldots,b^{R}$ for example $x$, conventional evaluation obtains a single point estimate $y$ of the ground truth through deterministic aggregation (top). Instead, we perform statistical aggregation by Bayesian inference (bottom), using an algorithm that samples plausibilities $\lambda$ that is probability vectors over conditions - from a posterior distribution $p(\lambda|b,x)$. As plausibilities are probability vectors, even a single sample captures inherent uncertainty; for example, a posterior plausibility sample allocating high probability mass to several conditions indicates higher uncertainty about the underlying case. In contrast, annotation uncertainty is represented by variation in plausibility samples. When no variation exists in the plausibility samples, the only remaining source of uncertainty is inherent uncertainty. We control variance using a single annotator reliability parameter, which the user can choose. This choice could be informed by a domain expert who designed the rating task or tuned using rating data (similar to crowd sourcing work [116, 121]). If selecting an appropriate rater reliability proves difficult, evaluating across a range of reliabilities is the best alternative.”
We propose measuring ground truth uncertainty and evaluating a trained model’s performance using such plausibility samples from the above posterior in a Monte Carlo fashion. Specifically, we define a case based measure, annotation certainty, as the fraction of times that, e.g., the top-1 label from plausibility samples agrees. This allows us to quantify annotation uncertainty on individual examples as well as on whole datasets. Moreover, we define uncertainty-adjusted variants of common classification metrics such as top-$k$ accuracy or average overlap [114, 111] in a similar fashion by averaging these metrics across plausibility samples. These metrics explicitly take ground truth uncertainty into account during evaluation. As a result, we avoid unfairly penalizing or rewarding the model’s predictions if the ground truth is deemed uncertain.
1.3 Application to skin condition classification
We apply our framework to skin condition classification from images in dermatology following the setting of [58]. Annotations are expressed as differential diagnoses, which we formalize as partial rankings that may exhibit ties and generally include only a small fraction of conditions. Due to the difficulty of diagnosis solely from images, we observe significant disparity among these differential diagnoses. In the literature, an established technique for aggregating these partial rankings is inverse rank normalization (IRN) [58]. This technique provides a single point estimate of the probabilities of the underlying conditions. Then, the condition attaining the highest IRN probability is chosen as the ground truth label. Evaluating e.g. the accuracy of an AI model in this way ignores any potential ground truth uncertainty. As a result, this evaluation is overconfident and ignores alternative scenarios where other conditions could have been more probable. This is particularly problematic in settings where AI models are meant to output multiple alternatives themselves (e.g., prediction sets or differential diagnosis). Moreover, this approach is unable to capture the impact that individual annotations may have on metrics or determine when improvements are statistically significant.
As part of our framework, we propose two approaches for probabilistic aggregation of differential diagnoses expressed as partial rankings: First, we consider a re-interpretation of IRN as a point estimator of a categorical distribution. Based on this interpretation, we propose a Probabilistic IRN (PrIRN) where we model the posterior distribution over plausibilities as a Dirichlet distribution. Second, we consider the Plackett–Luce (PL) model [75, 61] which models probability distributions over full rankings through sampling without replacement. We adapt PL to the case of partial rankings, overcoming several technical challenges. Based on these models, our analysis highlights that ignoring uncertainty leads to unrealistically confident and optimistic estimates of classifier performance. We observe that standard evaluation on curated labels that are derived from point estimates disregards large variations in performance, and hide the fact that results can be over-sensitive to the opinions of a single reader. We discussed these results with dermatologists and highlight medical implications such as the inability to clearly categorize cases by risk.
1.4 Summary of contributions
  1. 1.
    We address the problem of ground truth uncertainty in evaluating medical AI systems, comprising inherent uncertainty due to limited observational information and annotation uncertainty arising from finite and imperfect annotators.
  2. 2.
    We argue that conventional, deterministic approaches to aggregating annotations implicitly ignore this uncertainty. We propose a Bayesian framework that explicitly aggregates annotations into a posterior over plausibilities (distributions over classes).
  3. 3.
    We introduce annotation certainty to measure ground truth uncertainty on individual cases and entire datasets, and propose uncertainty-adjusted variants of classification metrics to account for this uncertainty during evaluation.
  4. 4.
    We demonstrate our framework on a skin condition classification task where annotations are expressed as differential diagnoses and aggregated using either probabilistic IRN (PrIRN) or Plackett-Luce (PL), showing that conventional evaluation overestimates performance and disregards significant variation due to ground truth uncertainty.
@@ 1.4 Summary of contributions
2 Evaluation with uncertain ground truth from differential diagnoses
@@
This section introduces our framework for evaluation of AI systems with ground truth uncertainty. As this paper focuses on a case study in dermatology, we particularly concentrate on its application to the differential diagnosis annotations from prior work [58]. We start by introducing notation and give an intuition of our approach on a toy example as well as a concrete example from our dermatology dataset. We then formalize this intuition and present two concrete statistical models we use for modeling and aggregating differential diagnosis annotations. Based on these models, we present measures for annotation uncertainty as well as uncertainty adjusted performance metrics for evaluating AI models.
@@ 2.1 Notation and introductory examples
Fig. 2: Illustration of annotation certainty on a toy dataset. Left: Synthetic dataset showing one-dimensional observations $x$ and corresponding true distribution $p(y|x)$ over labels $y$ for three imaginary labels Benign1, Benign2 and Cancer. Example (ii) is particularly ambiguous in the sense that $p(y|x)$ is not crisp. Middle left: For illustration, we plot the distributions $p(y|x)$ as “diamonds” on the 3-simplex. Middle and middle right: Modeling aggregation of annotator opinions $b$ statistically, see text, allows us to sample plausibilities $\lambda^{(m)}\sim p(\lambda|b,x)$. The spread of these plausibilities around the actual true distributions $p(y|x)$ captures the annotation uncertainty. This is influenced by the annotations as well as a prior reliability parameter $\gamma$ ($\gamma=1$ and $\gamma=100$ in this case). This parameter reflects our prior trust in the annotators. Right: Measuring how often the top-1 label changes given plausibility samples $\lambda^{(m)}$, we can compute annotation certainty for all examples. We plot annotation certainty across different reliabilities $\gamma$ (black to gray) indicating the effect of different prior trust levels.
For illustration and introducing notation, we consider the synthetic toy dataset of Figure 2 (left). Here, observations $x$ are illustrated as one-dimensional on the x-axis and we plot the true distribution $p(y|x)$ for three different classes (blue, orange and violet) on the y-axis. These could be different diagnoses, such as cancer and two benign conditions, benign1, benign2. For illustration, we pick three examples, (i), (ii), and (iii). These could correspond to patients to diagnose. Here, the latter two examples, (ii) and (iii), are inherently ambiguous: the corresponding distribution over classes $p(y|x)$ is not crisp as can be observed at the corresponding intersecting dotted line. For these cases, multiple classes have significant probability $p(y|x)$. For example, for example (ii), all three classes have roughly equal probability. We can visualize this more intuitively on a 3-simplex where the corners would correspond to crisp distributions (100% probability for cancer, benign1 or benign3), see the diamonds in Figure 2 (middle left). Of course, in practice, we never observe the true distributions $p(y|x)$ for these examples (represented as diamonds on the graph).
@@
To address this, we assume access to a finite set of annotations $b$ obtained from human expert raters. For simplicity, in this example we assume these annotations to be single labels, i.e., we assume that each rater samples a single label from the true $p(y|x)$ (or what they believe to be the true distribution). Then, we aggregate these opinions to obtain an approximation $\hat{\lambda}$ of $\lambda:=p(y|x)$. This can be achieved by counting frequencies. A ground truth label is then selected using, e.g., majority voting. This represents how labels for many common benchmarks such as CIFAR and ImageNet [85] have been obtained and this process is illustrated in Figure 1 (top). Here, we refer to $\hat{\lambda}$ as plausibilities as they construct a categorical distribution denoted $p(y|\hat{\lambda})$ over labels from which the $\arg\max$ corresponds to the majority voted label. Unfortunately, this approach ignores any uncertainty present in the annotations since the rich information in these plausibilities is discarded in favor of a single, majority-voted label.
@@
Instead, we intend to work directly with these plausibilities. To this end, we assume a distribution $p(\lambda|b,x)$ over plausibilities. For the toy example in Figure 2, we use a Dirichlet distribution with concentration parameters reflecting the annotator opinions as well as a prior reliability parameter ($\gamma$ in Figure 2). This reliability parameter will quantify our a priori trust in the annotators. For example, we expect reliability to increase with the number of annotators or their expertise and training. For now, we consider reliability to be a variable parameter that affords us various perspectives on the data. This is illustrated in Figure 2 (middle and middle right), showing plausibilities $\lambda^{(m)}\sim p(\lambda|b,x)$, $m\in[M]$, sampled from the statistical aggregation model on the 3-simplex. The spread in these plausibilities represents annotation uncertainty and can be reduced using a higher reliability. The position on the simplex represents inherent uncertainty which is lower for plausibilities close to the corners such as (i) and higher for plausibilities close to the center such as (ii). Then, we compute the top-1 label for every sampled $\lambda^{(m)}$, i.e., the label with the largest plausibility $\arg\max_{k\in[K]}\lambda^{(m)}_{k}$. This allows us to measure annotation certainty (defined formally in Section 2.4) – if the top-1 label is always the same label, there is no annotation uncertainty. Specifically we can measure the fraction of plausibilities $\lambda^{(m)}$ among all $M$ samples where the top-1 label is $k$. Taking the maximum over all labels $k\in[K]$ defines annotation certainty. In Figure 2 (right) we plot this annotation certainty for different reliabilities (black and gray) and we clearly see that annotation certainty is consistently low for example (ii), while annotation certainty increases with higher reliability for example (iii). This follows our approach as depicted in Figure 1 (bottom) and allows us to easily summarize annotation uncertainty across the whole dataset.
Input $x$
Annotations $b$
A0: {Pyogenic granuloma (Low)} {Hemangioma (Med)} {Melanoma (High)}
A1: {Angiokeratoma of skin (Low)} {Atypical Nevus (Med)}
A2: {Hemangioma (Med)} {Melanocytic Nevus (Low), Melanoma (High),
O/E - ecchymoses present (Low)}
A3: {Hemangioma (Med), Melanoma (High), Skin Tag (Low=}
A4: {Melanoma (High)}
A5: {Hemangioma (Med)} {Melanoma (High)} {Melanocytic Nevus (Low)}
*-4px
Fig. 3: An example input and corresponding expert dermatologist annotations from the dataset introduced in [58]. We highlight the conditions’ risk categorizations in parentheses and indicate ranks using curly brackets where tied conditions are grouped together. We further underline the majority voted label used for evaluation in [58], “Hemangioma”. For contrast, we mark an alternative, high-risk condition, “Melanoma”, in italics. There is clearly significant disagreement among experts, highlighting high ground truth uncertainty which could stem from either inherent or annotation uncertainty.
The main challenge in applying the aforementioned approach to our dermatology case study is that the expert annotations collected are not single labels per example. As illustrated in Figure 3, the annotations collected in [58] come in the form of differential diagnoses. Mathematically, differential diagnoses can be formalized as partial rankings, including a limited number of possible conditions that are ranked while allowing for ties (see Section 2.3). This has to be accounted for in the aggregation model $p(\lambda|b,x)$, which will get more complex compared to the simple Dirichlet model above. The example in Figure 3 highlights this difficulty on a concrete example with high level of disagreement among expert annotators. Here, the majority-voted label (used for evaluation in [58]; see Section 2.3 for formal definition) is “Hemangioma”, a medium-risk condition. However, this ignores a higher-risk, cancerous condition, “Melanoma” that is also present in many of the provided differential diagnoses. Clearly, limiting evaluation to “Hemangioma” is overly optimistic and disregards the rich information contained in the annotations. In our framework, we expect the top-1 label to switch frequently between “Hemangioma” and “Melanoma” for different samples $\lambda^{(m)}\sim p(\lambda|b,x)$, indicating high uncertainty. Adjusting $p(\lambda|b,x)$ to different settings makes this approach very flexible, see A for further examples.